# MARS: Optimizing Dual-System Deep Research via Multi-Agent Reinforcement Learning

### v1

@@ 3.5 Analysis of Multi-agent RL Process
(d) Usage Ratio of Python
(e) Usage Ratio of Google Search
(f) Usage Ratio of Google Scholar
(g) Min. Response Len. (Sys. 1)
(h) Mean Response Len.
(i) Max. Response Len. (Sys. 2)
Figure 3: Comprehensive analysis of our RL training process. The $x$-axis represents training steps for all subfigures. (a-c) Core performance metrics on HLE score (randomly select 320 questions), training reward, and tool usage frequency per question. (d-f) Evolution of tool selection preferences across three available tools. While Google Search emerges as the predominantly chosen tool due to our training data distribution, we maintain all tools to preserve System 2’s autonomous tool selection capability for diverse scenarios. (g-i) Response length distributions showing minimum (predominantly System 1), mean, and maximum (predominantly System 2) response lengths. Training was terminated after step 150 due to consistently exceeding our preset length constraints.
To gain deeper insights into our multi-agent, multi-turn, multi-tool RL process, we conduct a comprehensive analysis of MARS’s training dynamics. Figure 3 illustrates various aspects of the training progression, revealing several key patterns:
First, examining the core performance metrics, we observe a consistent and stable improvement in HLE score (Fig. 3(a)) from approximately 2% to over 10% on the HLE subset, demonstrating the effectiveness of our approach. This performance gain correlates with the training reward curve (Fig. 3(b)), which stabilizes around 0.4 after an initial increase. Interestingly, the number of tools used per question (Fig. 3(c)) also shows an upward trend, increasing from approximately 1 to over 2 by the end of training, indicating that MARS learns to leverage multiple tools for addressing complex questions. Second, the evolution of tool selection preferences (Fig. 3(d)-3(f)) reveals that while Google Search emerges as the predominantly chosen tool (reaching nearly 98% usage), our model maintains the capability to utilize Python and Google Scholar when appropriate. This balanced tool selection strategy allows System 2 to autonomously choose the most suitable tools for different question types, enhancing the model’s versatility across diverse reasoning scenarios. Third, the response length dynamics (Fig. 3(g)-3(i)) provide valuable insights into the operation of our dual-system architecture. The minimum response length (mostly System 1) and maximum response length (mostly System 2) both show increasing trends during training, suggesting that both systems learn to generate more comprehensive and informative outputs. The significant increase in mean response length indicates that MARS progressively develops more sophisticated reasoning capabilities as training advances.
3.6 Ablation Study
@@
To further investigate the impact of different external tools on reasoning performance, we conduct a comprehensive ablation study on the HLE benchmark, as shown in the Table 3.
We first observe that the model with all three tools achieves the best overall performance (7.38%), confirming that the combination of computational capabilities and diverse knowledge sources provides the most robust foundation for complex reasoning. Interestingly, the performance drop varies significantly depending on which tool is removed, revealing their differential importance across subject domains: removing Python affects Math (-1.54%) and Physics (-0.98%) most severely, while actually improving performance in Bio/Med (+0.46%) and Chemistry (+2.00%). Second, Google Search proves to be the most versatile single tool, with its removal causing the largest overall performance drop (1.38%), particularly in Physics (-2.95%). Then, Google Scholar contributes most significantly to CS/AI and Other categories, reflecting its value for domains with rapidly evolving research. These findings demonstrate that different tools contribute uniquely across subject domains. The complementary nature of these tools enables MARS to adapt its reasoning strategy based on the specific requirements of each question, highlighting the importance of our multi-tool approach in addressing the diverse challenges presented by complex reasoning tasks.
3.7 Information Volume Analysis
@@
In our MARS, each tool invocation supports multiple queries, with each query retrieving up to 10 complete web pages or 5 full academic papers, resulting in substantial information volume that System 1 effectively processes and distills for System 2’s reasoning. As shown in Table 4, MARS retrieves and processes a significant amount of information per question across all benchmarks. Particularly for HLE, the model accesses an average of 22.31 web pages and 0.17 academic papers per question, reflecting the complexity and knowledge-intensive nature of this benchmark. This analysis demonstrates that MARS effectively leverages a large volume of external information, highlighting the efficiency of our dual-system approach in processing and utilizing retrieved knowledge.
4 Related Work
Retrieval-Augmented Generation (RAG) has emerged as a crucial approach to overcome knowledge limitations of large language models by integrating external information sources [25, 35, 46, 61, 3, 21, 53, 60, 4, 29]. However, existing RAG systems typically struggle with either information overload when processing multiple lengthy documents (such as entire web pages or research papers) or loss of critical details when condensing information [10, 12, 48]. Our approach addresses these limitations through a dual-system framework where System 2 handles complex reasoning while System 1 efficiently processes retrieved information, enabling MARS to manage larger volumes of external information while maintaining reasoning quality across diverse knowledge-intensive tasks. Due to space constraints, additional related work is provided in Appendix A.
@@ 5 Conclusion
In this paper, we presented MARS, a multi-agent system for Deep Research that seamlessly integrates System 1’s fast, intuitive thinking with System 2’s deliberate reasoning. By creating a specialized division of labor between the two cognitive systems, MARS establishes an efficient collaborative framework where each component operates within its optimal domain. Furthermore, we proposed a multi-agent RL framework to simultaneously optimize both systems with bin-packing optimization and sample balancing strategies to enhance the collaborative efficiency. Extensive experiments demonstrate MARS’s superior performance and the effectiveness of our dual-system paradigm for complex reasoning tasks that require processing large volumes of external information.
@@ Appendix A Additional Related Work
Recent advances in language models have witnessed the emergence of LRMs, which specifically excel at deliberate, System 2-type thinking [20, 59, 15, 50, 2, 30, 65]. Despite these advances, current LRMs rely on static, parameterized knowledge acquired during pre-training, without access to external world information [4, 28, 22, 29]. Furthermore, when faced with simpler tasks, such as analyzing web pages or reading papers, LRMs may tend to overanalyze, consuming excessive reasoning tokens that limit their capacity to process large volumes of information. Our work addresses these limitations by integrating System 1’s fast, intuitive thinking with System 2’s deliberate reasoning in a multi-agent framework that leverages external knowledge sources.
@@
Multi-agent systems have gained significant attention in the LLM community as a means to tackle complex tasks through collaborative problem-solving [40, 54, 44, 16, 27, 4]. Recent work has explored various agent architectures, from simple role-playing approaches [26, 8] to more sophisticated frameworks with specialized agents handling different aspects of complex tasks [18, 52]. However, most existing multi-agent systems employ agents with similar cognitive architectures, typically all operating in a intuitive, System1-type thinking or deliberate, System 2-like reasoning mode. This homogeneity limits their ability to efficiently process large volumes of information while maintaining complex reasoning capabilities. Our approach diverges by implementing a heterogeneous multi-agent system where System 1 and System 2 agents possess complementary cognitive abilities, optimized through reinforcement learning [62, 64, 47] to collaboratively tackle complex research tasks requiring both efficient information processing and deliberate reasoning.
@@ Appendix B Training Data Filtering Pipeline
Given the absence of specialized training datasets for enhancing model capabilities required for challenging benchmarks such as HLE [42] (which demands web browsing, reasoning, and computation skills), we propose a data curation pipeline that selectively identifies suitable training examples from open-source datasets. As illustrated in Figure 4, our pipeline evaluates candidate examples based on three key dimensions: clarity of presentation, problem complexity, and solution correctness.
All candidate data are sourced from publicly available datasets, including MMIQC [31], WebInstructSub [63], GeneralThought [^2] https://huggingface.co/datasets/GeneralReasoning/GeneralThought-430K, and SuperGPQA [7]. To ensure quality for our dual-system training, we implement a rigorous multi-stage filtering process:
@@
    Clarity assessment: We employ an LLM (Qwen2.5-72B-Instruct [56]) to systematically evaluate the clarity of each prompt, filtering out those deemed ambiguous or poorly formulated, resulting in 99K high-quality prompts.
@@
    Challenge and Correctness Verification: To ensure both appropriate difficulty and answer verifiability, we perform a best-of-16 (BoN) sampling procedure using Qwen2.5-72B-Instruct with Google Search access. As shown in Figure 5, we plot the distribution of how many times (out of 16 attempts) the model correctly answers each question. We retain only questions with moderate difficulty (correctly answered 1-12 times out of 16 attempts), which serves two critical purposes: (1) eliminating trivial questions (answered correctly >12 times) and excessively difficult ones (never answered correctly), and (2) identifying questions with consistent, verifiable answers, as those never answered correctly may lack definitive solutions or contain ambiguous information.
@@
Figure 5: Distribution of Correct Number in Best-of-N ($N=16$). Questions answered correctly 1-12 times were retained for training, while those answered 0 times (potentially ambiguous or lacking definitive solutions) or >12 times (too trivial) were excluded.
To facilitate research advancement and reproducibility, we have open-sourced our curated 40K dataset. The detailed prompts used throughout our data curation pipeline are provided in Appendix D.2.
@@ C.1 Overall Algorithm
1: Question $q$, System 1 $\pi_{\text{sys}_{1}}$ and System 2 $\pi_{\text{sys}_{2}}$ in the same LLM, candidate tools, Maximum interaction turns $N_{\max}$.
2: $i\leftarrow 0$ $\triangleright$ Initialize interaction turns
3: $c_{0}\leftarrow\{q\}$ $\triangleright$ Initialize reasoning context for System 2
4: $c_{\text{sys}_{1}}\leftarrow\emptyset$ $\triangleright$ Initialize collection of System 1 input-output pairs
5: while $i<N_{\max}$ do
6:   $s_{i},(t_{i},p_{i})\leftarrow\pi_{\text{sys}_{2}}(c_{i})$ $\triangleright$ System 2 generates reasoning step $s_{i}$, (optional tool request $t_{i}$ and purpose $p_{i}$)
7:   if $t_{i}$ and $p_{i}$ are not Empty then $\triangleright$ Tool request present
8:    $\{o_{t_{i}}^{(1)},o_{t_{i}}^{(2)},\ldots,o_{t_{i}}^{(n_{t_{i}})}\}\leftarrow\text{Tool Call}(t_{i})$ $\triangleright$ Execute tool call
9:    $\{b_{1},\ldots,b_{m}\}\leftarrow\text{Bin-Packing}(o_{t_{i}}^{(1)},\ldots,o_{t_{i}}^{(n_{t_{i}})})$ $\triangleright$ $m$ chunks
10:    $\{d_{1},\ldots,d_{m}\}\leftarrow\{\pi_{\text{sys}_{1}}(b_{1}),\ldots,\pi_{\text{sys}_{1}}(b_{m})\}$ $\triangleright$ Parallel distillation (System 1)
11:    $c_{\text{sys}_{1}}\leftarrow c_{\text{sys}_{1}}\cup\{(b_{j},d_{j})\}_{j=1}^{m}$ $\triangleright$ Collect all input-output pairs
12:    $\tilde{o}_{t_{i}}\leftarrow\cup_{j=1}^{m}d_{j}$ $\triangleright$ Combine all distilled information
13:    $c_{i+1}\leftarrow c_{i}\cup\{s_{i},t_{i},p_{i},\tilde{o}_{t_{i}}\}$ $\triangleright$ Update System 2’s context with tool results
14:   else
15:    $c_{i+1}\leftarrow c_{i}\cup\{s_{i}\}$ $\triangleright$ Update context with reasoning only
16:    if <answer> </answer> in $s_{i}$ then $\triangleright$ Check if answer is provided
17:      break
18:   $i\leftarrow i+1$
19:return Rollout trajectory $c_{i}$ and System 1 input-output pairs $c_{\text{sys}_{1}}$

### v2

@@ 4.1 Evaluation Datasets and Metrics
Figure 3: Analysis of our RL training process. The $x$-axis represents training steps. (a) HLE score progression (evaluated on 320 random questions) shows consistent improvement from $\sim$2% to over 10%. (b) Training reward stabilizes around 0.4 after initial increase. (c) Tool usage per question increases from $\sim$1 to over 2, indicating that MARS learns to leverage multiple tools for complex questions. Extended analysis including tool selection preferences and response length dynamics is provided in Appendix D.1.
4.2 Baselines and Implementation Details
To evaluate the effectiveness of MARS, we compare our method with the following baselines. (1) Direct Reasoning: Models that directly answer questions without external knowledge, including open-source models (Qwen2.5 series (Yang et al., 2024) and QwQ-32B (Team, 2024)) and powerful proprietary models (DeepSeek-R1-671B (Guo et al., 2025), GPT-4o (Hurst et al., 2024), o1 (OpenAI, 2024), Claude 3.7 Sonnet (Anthropic, 2025), Gemini 2.5 Pro (Google, 2025b), o3(high) (OpenAI, 2025c)), and o4-mini(high) (OpenAI, 2025b). (2) Advanced RAG Reasoning Methods: We consider the standard RAG that retrieve top-10 documents based on the question, and several iterative RAG methods, including Self-RAG (Asai et al., 2024), InstructRAG (Wei et al., 2025), Auto-RAG (Yu et al., 2024b) and C-3PO (Chen et al., 2025a). (3) R1-like Reasoning with Search: Methods that integrate external knowledge into R1-like reasoning, including Search-o1 (Li et al., 2025a), Search-R1 (Jin et al., 2025), WebThinker (Li et al., 2025b), and OpenAI Deep Research (OpenAI, 2025a).
Implementation and Fair Comparison. We initialize our policy model with Qwen2.5-7B-Instruct (Yang et al., 2024) and Qwen3-8B (Yang et al., 2025) and design dedicated prompts for System 1 and System 2 (Appendix E) to facilitate multi-agent reinforcement learning. Crucially, MARS operates under a “Zero RL” setting: we train directly from the base model without any Supervised Fine-Tuning (SFT) or policy distillation from stronger models. This contrasts with baselines such as WebThinker, which employs SFT on curated trajectories before RL training. Table 1 summarizes the key differences in training configurations. Despite using smaller models (7B/8B vs. 32B/72B) and no SFT warmup, MARS achieves superior performance, validating the effectiveness of our multi-agent RL framework rather than relying on model scale or supervised data. All training data are curated from public sources (Appendix B), ensuring data source parity with baselines. Additional implementation details are provided in Appendix C.
4.3 Main Results on Humanity’s Last Exam
Table 2 compares MARS with various baselines on HLE, a challenging benchmark requiring sophisticated reasoning and up-to-date knowledge across multiple disciplines.
MARS achieves 7.38% accuracy (8.17% with Qwen3-8B), a 3.86% improvement over the base model, outperforming larger models including WebThinker (32B) and C-3PO (72B). The performance gap with proprietary models like Claude 3.7 Sonnet (7.89%) and o1 (7.75%) is notably small, especially considering MARS uses only 7B/8B parameters. Beyond the overall score, Table 2 suggests that our gains are broad rather than isolated to a single subject area, consistent with the goal of building a general deep-research agent. We attribute these improvements to the complementary roles of the two systems: System 2 concentrates its long-form reasoning budget on planning and synthesis, while System 1 absorbs the high-throughput reading burden by filtering and compressing retrieved content. Importantly, this separation mitigates a common failure mode of tool-augmented reasoning—context dilution. Instead of appending raw documents into the reasoning context, System 1 produces purpose-conditioned evidence snippets, which makes subsequent reasoning steps more stable and less sensitive to retrieval noise.
4.4 Main Results on Knowledge-intensive Reasoning
Table 3 presents results on 7 knowledge-intensive QA tasks covering both single-hop and multi-hop reasoning. MARS consistently outperforms all baselines, achieving 8.95% average improvement over C-3PO. The gains are particularly pronounced on multi-hop tasks (+12.2% over C-3PO), demonstrating that our framework excels at complex reasoning chains that require multiple retrieval–reasoning iterations. These results highlight that the benefit of co-evolution is not limited to HLE. On single-hop QA, our method improves factual recall by using lightweight, purpose-guided retrieval and distillation, reducing hallucinations caused by relying on stale parametric knowledge. On multi-hop QA, the advantage is larger because errors compound across steps: missing one entity or relation early can derail the entire chain.
We attribute the improvements to two complementary behaviors learned during joint training. First, System 2 learns to decompose questions into targeted sub-queries and to interleave tool use with reasoning, maintaining a coherent reasoning trace instead of attempting to “read everything” in-context. Second, System 1 learns purpose-conditioned distillation that preserves entities, relations, and intermediate constraints (rather than producing generic summaries), which is critical for multi-hop settings where fine-grained details determine whether subsequent hops are correct. Overall, these trends suggest that MARS improves both evidence acquisition (better tool-use decisions) and evidence utilization (more faithful, task-relevant distillation), leading to robust gains across diverse knowledge-intensive benchmarks.
4.5 Analysis of Multi-agent RL Process
Figure 3 illustrates key aspects of our multi-agent RL training dynamics. We observe consistent improvement in HLE score from approximately 2% to over 10%, correlating with reward stabilization around 0.4. Meanwhile, the average number of tools used per question increases from $\sim$1 to over 2, suggesting a shift from “single-shot” retrieval toward iterative evidence gathering.
@@ 4.5 Analysis of Multi-agent RL Process
Table 5: Ablation on System 1 co-evolution (HLE).
Method Bio/Med Chem. CS/AI Eng. Hum. Math Phys. Other Avg.
Qwen2.5-7B (Base) 5.42 3.00 1.76 3.22 4.66 3.58 1.98 4.00 3.52
MARS (Full) 12.66 3.00 5.75 4.83 11.92 6.46 6.43 7.42 7.38
     w/o System 1 9.95 5.00 3.98 6.45 7.77 4.92 3.46 4.57 5.47
@@
Beyond the overall upward trend, Figure 3 reveals two qualitative phases. In early training, the model is conservative in tool use and often under-explores external evidence, leading to low reward and slow score improvement. As training proceeds, the policy learns to invoke tools more frequently on harder questions and to sustain longer interaction trajectories; this coincides with higher reward and improved accuracy. Notably, these trends are aligned: as training progresses, the model not only becomes better at deciding when to use tools, but also improves how to use them (i.e., decomposing a question into multiple targeted queries and integrating the resulting evidence). This is consistent with our design goal that System 2 should specialize in planning/tool-use, while System 1 specializes in high-throughput information processing. We find this behavior important because naive tool-augmented reasoning often fails due to noisy retrieval: issuing more tool calls can easily increase irrelevant context and distract reasoning. In MARS, System 1 mitigates this risk by purpose-conditioned distillation, while the balanced sampling and decoupled-gradient design ensures that System 1 is directly optimized for utility to System 2 rather than for generic summarization. Extended analysis of RL dynamics (tool preferences and response-length trends) is provided in Appendix D.1.
4.6 Ablation Study
Tool Ablation on HLE. Table 4 shows that using all three tools achieves the best performance (7.38%), highlighting that complex questions often require both reliable retrieval (Search/Scholar) and precise computation (Python). Several patterns are worth noting. (1) Google Search is the strongest single tool: removing Search yields the largest overall drop, consistent with Search providing broad, high-recall coverage across domains. (2) Python is domain-specialized: its removal hurts Math/Physics most, indicating that many questions require exact arithmetic, symbolic manipulation, or unit-aware calculations that are brittle under purely textual reasoning. (3) Scholar complements Search: although Scholar is used less frequently, it benefits research-heavy categories (e.g., CS/AI, Other) where reliable answers may depend on technical terminology and up-to-date academic evidence. A key takeaway is that tools are not simply additive: removing a tool can sometimes improve a subset of categories, reflecting that the model must learn both tool selection and tool orchestration. This can happen when a weaker tool policy introduces noisy evidence or distracts the reasoning process. Our joint training setup mitigates this by encouraging System 2 to learn when a tool is worth the latency/cost and by training System 1 to extract only task-relevant evidence from tool outputs, reducing the chance that retrieval noise dilutes System 2’s context.
Information Volume. Table 6 shows that MARS processes substantial external information per question (e.g., many full web pages/papers), motivating the need for System 1’s efficient distillation. This analysis highlights a practical constraint for deep research: naively concatenating retrieved documents quickly saturates the context window and degrades reasoning due to irrelevant or redundant text. Our bin-packing + System 1 distillation pipeline makes the information volume tractable by converting many long documents into a small set of evidence-focused chunks that System 2 can reliably consume.
4.7 Ablation on System 1 Co-evolution
[…]
@@ 5 Conclusion
We presented MARS, a co-evolution framework that jointly trains System 1 (efficient information processing) and System 2 (deliberate reasoning) via multi-agent RL, with decoupled gradients, bin-packing, and advantage-balanced sampling to enable effective collaboration. Under a Zero RL setting (no SFT), MARS (8B) achieves 8.17% on HLE and improves performance by 8.9% on average across 7 knowledge-intensive QA tasks; ablations (e.g., removing System 1) confirm that the gains come from learned co-evolution rather than tool access alone.
Impact Statement
Our work improves the reliability of deep research by enabling language models to efficiently synthesize up-to-date information from external tools while maintaining strong multi-step reasoning. This capability may benefit knowledge-intensive applications such as education and scientific literature review. Potential risks include misuse for scalable misinformation or low-quality automated reports, and errors introduced by noisy retrieval or imperfect tool outputs. We therefore emphasize transparent evaluation, careful deployment with human oversight in high-stakes settings, and continued research on robustness and provenance-aware information distillation.
@@ Appendix A Additional Related Work
Recent advances in language models have witnessed the emergence of LRMs, which specifically excel at deliberate, System 2-type thinking (Jaech et al., 2024; Yu et al., 2024a; Guo et al., 2025; Team, 2024; Anthropic, 2025; Li et al., 2025c; Ziabari et al., 2025; Chen et al., 2025c). Despite these advances, current LRMs rely on static, parameterized knowledge acquired during pre-training, without access to external world information (Chen et al., 2025a; Li et al., 2025a; Jin et al., 2025; Li et al., 2025b). Furthermore, when faced with simpler tasks, such as analyzing web pages or reading papers, LRMs may tend to overanalyze, consuming excessive reasoning tokens that limit their capacity to process large volumes of information. Our work addresses these limitations by integrating System 1’s fast, intuitive thinking with System 2’s deliberate reasoning in a multi-agent framework that leverages external knowledge sources.
Reasoning Models with External Tools.
Recent work has explored integrating external tools into reasoning models (Chen et al., 2024a, b; Li et al., 2025a; Jin et al., 2025; Li et al., 2025b; Chen et al., 2025a, b). Search-R1 (Jin et al., 2025) and Search-o1 (Li et al., 2025a) directly feed raw search results to reasoning models, which can overwhelm the context with irrelevant information. WebThinker (Li et al., 2025b) relies on a fixed extraction prompt to process retrieved documents before reasoning, while C-3PO (Chen et al., 2025a) uses supervised fine-tuning with curated trajectories but does not co-evolve the information processor and reasoner. In contrast, MARS enables co-evolution between System 1 and System 2 through multi-agent RL, allowing System 1 to learn task-specific information distillation that adapts to System 2’s needs.
@@
Multi-agent systems have gained significant attention in the LLM community as a means to tackle complex tasks through collaborative problem-solving (Park et al., 2023; Wu et al., 2023; Li et al., 2023; Hong et al., 2023; Qian et al., 2023; Han et al., 2024; Li et al., 2024; Chen et al., 2025a). Recent work has explored various agent architectures, from simple role-playing approaches (Li et al., 2023; Du et al., 2023) to more sophisticated frameworks with specialized agents handling different aspects of complex tasks (Hong et al., 2023; Wang et al., 2023). However, most existing multi-agent systems employ agents with similar cognitive architectures, typically all operating in a intuitive, System1-type thinking or deliberate, System 2-like reasoning mode. This homogeneity limits their ability to efficiently process large volumes of information while maintaining complex reasoning capabilities. Our approach diverges by implementing a heterogeneous multi-agent system where System 1 and System 2 agents possess complementary cognitive abilities, optimized through reinforcement learning (Yuan et al., 2023; Zhu et al., 2024; Shao et al., 2024) to collaboratively tackle complex research tasks requiring both efficient information processing and deliberate reasoning.
@@ Appendix B Training Data Filtering Pipeline
Given the absence of specialized training datasets for enhancing model capabilities required for challenging benchmarks such as HLE (Phan et al., 2025) (which demands web browsing, reasoning, and computation skills), we propose a data curation pipeline that selectively identifies suitable training examples from open-source datasets. As illustrated in Figure 4, our pipeline evaluates candidate examples based on three key dimensions: clarity of presentation, problem complexity, and solution correctness.
All candidate data are sourced from publicly available datasets, including MMIQC (Liu et al., 2025), WebInstructSub (Yue et al., 2024), GeneralThought [^2] https://huggingface.co/datasets/GeneralReasoning/GeneralThought-430K, and SuperGPQA (Du et al., 2025). To ensure quality for our dual-system training, we implement a rigorous multi-stage filtering process:
@@
    Clarity assessment: We employ an LLM (Qwen2.5-72B-Instruct (Yang et al., 2024)) to systematically evaluate the clarity of each prompt, filtering out those deemed ambiguous or poorly formulated, resulting in 99K high-quality prompts.
@@
    Challenge and Correctness Verification: To ensure both appropriate difficulty and answer verifiability, we perform a best-of-16 (BoN) sampling procedure using Qwen2.5-72B-Instruct with Google Search access. As shown in Figure 5, we plot the distribution of how many times (out of 16 attempts) the model correctly answers each question. We retain only questions with moderate difficulty (correctly answered 1-12 times out of 16 attempts), which serves two critical purposes: (1) eliminating trivial questions (answered correctly ¿12 times) and excessively difficult ones (never answered correctly), and (2) identifying questions with consistent, verifiable answers, as those never answered correctly may lack definitive solutions or contain ambiguous information.
@@
Figure 5: Distribution of Correct Number in Best-of-N ($N=16$). Questions answered correctly 1-12 times were retained for training, while those answered 0 times (potentially ambiguous or lacking definitive solutions) or ¿12 times (too trivial) were excluded.
To facilitate research advancement and reproducibility, we have open-sourced our curated 40K dataset. The detailed prompts used throughout our data curation pipeline are provided in Appendix E.2.
@@ C.1 Overall Algorithm
0: Question $q$, System 1 $\pi_{\text{sys}_{1}}$ and System 2 $\pi_{\text{sys}_{2}}$ in the same LLM, candidate tools, Maximum interaction turns $N_{\max}$.
[…]
4:while $i<N_{\max}$ do
5:  $s_{i},(t_{i},p_{i})\leftarrow\pi_{\text{sys}_{2}}(c_{i})$ {System 2 generates reasoning step $s_{i}$, optional tool request $t_{i}$ and purpose $p_{i}$}
6:  if $t_{i}$ and $p_{i}$ are not Empty then
7:   $\{o_{t_{i}}^{(1)},o_{t_{i}}^{(2)},\ldots,o_{t_{i}}^{(n_{t_{i}})}\}\leftarrow\text{Tool Call}(t_{i})$ {Execute tool call}
8:   $\{b_{1},\ldots,b_{m}\}\leftarrow\text{Bin-Packing}(o_{t_{i}}^{(1)},\ldots,o_{t_{i}}^{(n_{t_{i}})})$ {$m$ chunks}
9:   $\{d_{1},\ldots,d_{m}\}\leftarrow\{\pi_{\text{sys}_{1}}(b_{1}),\ldots,\pi_{\text{sys}_{1}}(b_{m})\}$ {Parallel distillation}
10:   $c_{\text{sys}_{1}}\leftarrow c_{\text{sys}_{1}}\cup\{(b_{j},d_{j})\}_{j=1}^{m}$ {Collect input-output pairs}
11:   $\tilde{o}_{t_{i}}\leftarrow\cup_{j=1}^{m}d_{j}$ {Combine distilled information}
12:   $c_{i+1}\leftarrow c_{i}\cup\{s_{i},t_{i},p_{i},\tilde{o}_{t_{i}}\}$ {Update context with tool results}
13:  else
14:   $c_{i+1}\leftarrow c_{i}\cup\{s_{i}\}$ {Update context with reasoning only}
15:   if answer tag detected in $s_{i}$ then
16:    break {Answer provided}
17:   end if
18:  end if
19:  $i\leftarrow i+1$
20:end while
21:return Rollout trajectory $c_{i}$ and System 1 input-output pairs $c_{\text{sys}_{1}}$